OpenStudy (anonymous):

I have a question: So! A box has square base and total surface area equal to 12 m2. (a) Express its volume as a function of x, the length of one side of the base. (b) Find the maximum volume of such a box. I am not sure how to set up this problem. Do I need to use the volume formula of a box or something to set the equation as a function of x? Thank you!

OpenStudy (anonymous):

volume is the area of the base times the height. the area of the base is \(x^2\) since it is a square your job is to find an expression for the length of the height in terms of \(x\) to do that note the surface area is \(2x^2+4xh=12\) solve this for \(h\) in terms of \(x\)

OpenStudy (anonymous):

you get \[4xh=12-2x^2\] \[h=\frac{12-x^2}{4x}=\frac{3}{x}-\frac{x}{4}\]

OpenStudy (anonymous):

you get \[4xh=12-2x^2\] \[h=\frac{12-x^2}{4x}=\frac{3}{x}-\frac{x}{4}\]

OpenStudy (anonymous):

question!!!

OpenStudy (anonymous):

multiply by \(x^2\) to get \(V(x)\) then find the max

OpenStudy (anonymous):

how did you get the 2x^2 and the 4xh?

OpenStudy (anonymous):

Sorry, I am kind of slow at this.

OpenStudy (anonymous):

Oh and what happened to the 2 when you were putting the h by itself? shouldn't it be 12-2x^2/4x?

OpenStudy (anonymous):

i drew a picture you have a box top and bottom have area \(x^2\) and then 4 sides have area \(xh\)